Erik is 32 years younger than Megan. 3 years ago, Megan's age was 3 times Erik's age. How old is Erik now?
Found 3 solutions by ikleyn, josgarithmetic, MathTherapy:
Answer by ikleyn(52790)

.

Three years ago, the difference of their ages was 32 years, and the older was 3 times as old as the younger.


It means that the difference of 32 years is two times the age of the younger three years ago.


So, 3 years ago Eric was 16 years old;  Megan was 16 + 32 = 48 years old.


Hence, now Eric is 19 years old;  Megan is 19 + 32 = 51 years old.    ANSWER


Answer by josgarithmetic(39618)
Answer by MathTherapy(10552)


Erik is 32 years younger than Megan. 3 years ago, Megan's age was 3 times Erik's age. How old is Erik now?
Let Erik's age be E
Them Megan's is: = E + 32
We then get: E + 32 - 3 = 3(E - 3)
E + 29 = 3E - 9
E - 3E = - 9 - 29
- 2E = - 38
Erik, or